On July 2, 2019, you decided to start up a new business – SmartReader Inc., an off-campus bookstore where students can buy textbooks and supplies at reduced prices. The following are summary transactions for the period July 2, 2019 to October 31, 2019, the company’s year end.

  1. You and several other shareholders invested $20,000 in return for shares in the company.
  2. A suitable location is found and rent is $1,000 per month. The lease agreement is for one year (i.e., ends on June 30, 2020). The entire rent for one year is due upon signing of the lease on July 2, 2019.
  3. A 6-month bank loan in the amount of $20,000 was obtained on August 1, 2019. The annual interest rate is 9%.
  4. Furniture and fixtures are purchased at a cost of $15,000. These are purchased for cash.
  5. Books and supplies of $50,000 was purchased on account.
  6. An insurance policy was purchased for $1,200 cash. The policy takes effect on July 2, 2019 and expires on June 30, 2020.
  7. An additional $120,000 of inventory was purchased on account.
  8. Sales for the period ended October 31, 2019 were:

Cash sales - $190,000

Sales on account - $6,000

       Cost of goods sold related to the sales amounts to $130,000.

  1. A total of $4,000 of the sales made on account were collected.
  2. Additional cash disbursements for the year were as follows:

Wages and salaries

$36,000

Advertising

2,000

Miscellaneous expense

1,500

Dividends to shareholders

10,000

Payments on account re: purchases of inventory

120,000

$169,500

The following adjustments at year end must be made:

  1. The furniture and fixtures are expected to last a total of 10 years with no residual value. The straight-line method is used.
  2. The adjustment for insurance expense.
  3. The interest payable on the bank loan. Credit to “Accrued Liabilities”.
  4. Invoices received but not yet paid amount to $700 for miscellaneous expenses. Credit Accrued Liabilities.
  5. Employees are owed a total of $600 for wages and salaries. Credit to “Accrued Liabilities”.
  6. The adjustment for rent expense.
  7. The expected income tax rate is 30%. Credit to “Accrued Liabilities”.

Required: Prepare the following statements with proper headings:

  • Multi-step income statement
  • Statement of Changes in Equity
  • Statement of Financial Position
Answer #1
SmartReader Inc.
Income Statement for year ending October 31,2019
Particulars $ $
Revenue
Cash sales 190000
Sales on account 6000
Total Revenue 196000
Expenses
Cost of goods sold 130000
Wages and salaries 36600 =36000+600
Advertising 2000
Miscellaneous expense 2200 =1500+700
Depreciation on furniture and fixtures 500 =15000/10*4/12
Insurance Expense 400 =1200*4/12
Interest expense on the bank loan 450 =20000*9%*3/12
Rent Expense 4000 =1000*4
Total Expenses 176150
Income before Taxes 19850
-Tax @30% -5955
Net Income 13895
-Dividends to shareholders -10000
Transferred to Retained Earnings 3895
SmartReader Inc.
Statement of Financial Position as at October 31,2019
Assets $ $
Non-Current Assets:
Furniture and Fixtures 14500 14500 =15000-500
Current Assets:
Prepaid Insurance 800 =1200-400
Inventory 40000 =50000+120000-130000
Accounts Receivable 2000 =6000-4000
Cash 48300 91100 =20000+20000-15000-1200+4000+190000-36000-2000-1500-10000-120000
Total Assets 105600
Equity & Liabilities $ $
Share Capital 20000
Retained Earnings 3895
Total Equity 23895
Non-Current Liabilities:
Bank Loan 20000 20000
Current Liabilities:
Accounts Payable 50000 =50000+120000-120000
Accrued Liabilities 7705 =450+700+600+5955
Rent Payable 4000 61705
Total Equity & Liabilities 105600
SmartReader Inc.
Statement of changes in equity for year ending October 31,2019
Share Capital Retained Earnings Total
Balance as on July 2, 2019 $0 $0 $0
Shares issued $20,000 $0 $20,000
Net Income $0 $13,895 $13,895
Less: Dividends $0 -$10,000 -$10,000
Balance as at October 31,2019 $20,000 $3,895 $23,895